OpenStudy (anonymous):

Fractions with algebra help?? Solve for x: the quantity of x plus 4 all over 2 = 7 x = 3 x = 5 x = 6 x = 10

OpenStudy (anonymous):

\[\frac{x+4}{2}=7\iff x+4=14\] is a start

OpenStudy (anonymous):

So x=10??

OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4 x = 15 x = 15 over 2 x = 9 x = 11 What about this one? @satellite73 @KamiBug

OpenStudy (anonymous):

\[\frac{1}{4}(2x-14)=4\] \[2x-14=16\] etc

OpenStudy (anonymous):

add 14 divide by 2

OpenStudy (anonymous):